A freeway segment with largely level terrain has an observed free-flow speed of approximately 70 mph. There are three lanes in each direction, a 3-ft lateral clearance, and approximately 1 interchange per mile. You collect some data and observe a volume of 3080 veh/h with a corresponding PHF = 0.88. You counted 154 trucks and buses and no RVs. From the driver behavior and what you already know about the area, you assume the drivers are all commuters.

What is the LOS for this freeway?
Answer #1

Step 1:

Find the heavy vehicle %

%=154x100/3080

%=0.05x100=5%

fhv=1/1+Pt(Et-1)

fhv=1/1+0.05(1.5-1)

fhv=0.975

Step 2:

Vp=V/(PHFxNxfhv)

Vp=3080/(3x0.88x0.975)

Vp=1190 pc/hr/ln

Step 3:

S=BFFS-adjustments for number of lanes, lateral clearance, interchanges

S=70-1.2-3-2.5

S=63.3 mph

D=Vp/S

D=1190/63.3

D=18.79 pc/mi/ln

The LOS for the given freeway is C